Subject: 3.5-m Users Committee minutes 11/19/07
From: Bruce Gillespie
Submitted: Thu, 22 Nov 2007 13:49:58 -0700
Message number: 1073
(previous: 1072,
next: 1074
up: Index)
APO 3.5-m Users Committee Phonecon, 11/19/07
Attending: Suzanne Hawley, Bruce Gillespie, Mark Klaene, Scott
Anderson, Michael Strauss, Remy Indebetouw, John Bally, Bill
Ketzeback, Jon Holtzman, Jon Fulbright
Absent: Al Harper
Minutes taken by Bruce Gillespie
**********************************
User feedback, comments from institutional representatives:
Princeton (Strauss) - Michael reported that everything has been fine
with observing. He added that there doesn't seem to be much of a
groundswell of support among Princeton users for the proposed wide-
field NIR imager, as several felt that NEWFIRM at KPNO would offer
more capability (a wider field). However, if multi-object
spectroscopy were available on the new instrument, that would make it
attractive.
UVa (Indebetouw) - Remy said that UVa was very happy with their recent
student field trip to APO, and that they got some good data. He
hasn't received much feedback regarding the proposed wide-field NIR
imager.
UW (Anderson, Hawley) - Scott had nothing to report from the user
side. Of the six people who have expressed opinions on the proposed
wide-field NIR imager, two were positive, three neutral, and one
negative.
CU (Bally) - John said that several CU astronomers had recently used
NIC-FPS (see report below). His users also are lukewarm about the
proposed wide-field NIR imager at APO, both because of NEWFIRM (see
above) and the WFCAM instrument at UKIRT. John added that he recently
hosted a successful CU student field trip at APO.
JHU (Fulbright) - Jon had nothing new to report, and had already given
JHU feedback on the proposed NIR imager.
NMSU (Holtzman) - Jon had no user feedback about recent observing.
Several faculty have given him positive votes for the proposed wide-
field NIR imager informally, but nothing in writing so it is hard to
judge what the NMSU sentiments are.
UChicago (Harper) - No report.
**********************************
Proposed wide-field NIR instrument followup - Hawley
Suzanne said that she and Bruce had spoken with the PI since our last
meeting. Since it now appears that there is only modest sentiment of
interest among our users for such an instrument at APO, we will
entertain an approach of hosting it at APO as a guest visiting
instrument, with only a modest level of support.
**********************************
Annual ARC Board of Governors meeting - Hawley
Bruce will send the Users Committee members the link to download the
3.5-m and Site annual reports. Suzanne said that the telescope report
featured the several student training sessions that have been held at
the site this year, science highlights, improved down-time statistics,
and instrument and telescope engineering accomplishments and plans.
The Board approved the 2008 budget request with a 5% overall increase
that includes augmentations to both the CIF projects and Sinking Fund
budgets. The rebuilding of the tertiary mirror mount was highlighted
as the next big CIF project. Suzanne said the Board was told of the
potentially negative impact to 3.5-m operations if there were no SDSS-
III project, and the Board members were encouraged to talk to their
institutions about joining.
**********************************
Discussion of telescope/instruments report - Klaene:
The report is given below. Mark walked through the highlights, and
there was some discussion:
Mark spoke of the recent good weather we have enjoyed at APO, although
we have had some nights with easterly winds (=> mediocre seeing).
Forecasters are predicting a warm and dry winter, which is good for
observing but bad for increased risk of spring forest fires. The
recent problems with tertiary mirror rotation may have been traced to
a clutch problem, which is being worked on. Telescope pointing has
been adjusted, and the problem with losing communications with SPIcam
(see below) illustrates the relative fragility of this instrument vis-
a-vis spare parts and documentation. Recall that the "P" in >10-year
old SPIcam stands for "Prototype".
*************************
3.5-m Telescope, Instruments, and CIF Projects Highlights, 10/11/07
through 11/14/07
Mark Klaene
0) Overview
Warm and dry weather dominated this past month with nearly no time
lost due to weather. Long-range forecast calls for a La Niņa year
with dryer and warmer-than-average weather over the winter. SDSS
construction continues without impacting day-to-day operations.
Several student groups from ARC institutions visited and received
training during this period.
1) Telescope
Tertiary rotation/clamp issues are still being investigated for a
single source problem; until this is resolved, users may see
occasional instances of tertiary rotation misbehaviour where small
amounts of observing time may be lost. Pointing has returned to
normal performance--one tracking problem was traced to slippage in the
drive box and corrected with a minor adjustment.
2) Instruments
o DIS: no problems this period
o SPIcam: fiber interface locked up but came back to life, spare
ordered
o NIC-FPS: 16-channel readouts are working, but the etalon is not
commissioned yet. Current issues: noise level higher than expected,
Fowler sampling is currently limited to 8, ion vacuum pump is running
high
o Echelle: Fritz Stauffer fixed the software problem with exposure
time stamps
o CorMASS: no problems this period
o Agile: optics for the focal reducer are getting AR coatings
o GFP: operational
o APOLLO: working well, including some remote operations
3) CIF projects
o Axis controller upgrades: initial testing yielded good results,
working on servo tuning and more sky testing during next engineering
time
o New rotator motor: modifications to mount were completed, waiting
for new axis controller
o Auto-focus: resumed work on software actor rewrite
o Tertiary: concept review completed
o Agile rotation: continued work on rotator and mount for TR2 port
**********************************
NIC-FPS status - Bally:
After noting problems with bias drifts during engineering tests, Corey
Wood (CU) found and fixed a bug in the instrument code that handles
the new 16-channel readouts. John feels that the camera side of NIC-
FPS is relatively stable, although there is some excess noise in the
frames that may be due to inadequate shielding of the fan-out board.
Consideration is being given to making a new fan-out board and
installing it next summer. Also, the ion pump is showing its age and
we're hoping to limp along with it until next summer--changing the fan-
out board and ion pump requires opening the dewar. Fowler sampling up
to F=15 is now supported by the instrument and control software, but
the current version of TUI restricts Fowler sampling to a limit of
F=8; it is recommended that all NIC-FPS frames use F=8 for now.
Upcoming on-sky engineering tests will focus on etalon issues. The
reference capacitor for one of the three etalon control axes probably
needs to be adjusted. More engineering tests are planned in Q1 '08.
After the upcoming tests are done, John will send updated user
information to the community. As of now, the etalon is not ready for
use, and users planning on future use of the etalon side of NIC-FPS
should contact John Bally for its current status.
**********************************
User documentation - Hawley/Holtzman:
An e-mail was recently sent to apo35-general that described the new
user information interface--Jon would be happy to receive additional
user comments. Michael said he had sent some comments, and had one
more coming. Suzanne pointed out that the new pages use a uniform
style, and contain up-to-date information. We need more information
that characterize the instruments, e.g., throughput, etc. Jon would
like to hear if there are missing items that users need to know
about. Suzanne said that there needs to be a tutorial on Fowler
sampling.
ACTION: Jon Bally will write a section with figures on Fowler
sampling for inclusion in the NIC-FPS user documentation.
**********************************
Echelle timing issue - Hawley/Klaene:
Mark reported that users have recently discovered that the time stamps
in the echelle data headers have been wrong, by as much as 10 minutes
or more. Fritz Stauffer has since put in a temporary patch that
monitors the time differences between the instrument and system
computers. A more permanent fix requires a planned computer upgrade.
If the past header time discrepancies are important to users, the
correct time of observations can be recovered from archived telemetry;
contact the site staff for assistance.
**********************************
Miscellaneous:
Bruce reported that we collect each year's User Committee minutes into
a single document for distribution at the annual Board of Governors
meeting. If anybody would like to have an electronic copy of these
compendia, please contact him.
Suzanne said that she is working on arrangement to experimentally
share a small number of nights between the ARC 3.5-m and WIYN
telescopes. As a result, during the 2nd half of 2008 we should have
access to ~3 nights of WIYN time. If interested, users should contact
Suzanne.
**********************************
Actions from earlier meetings - group
[open from previous months]:
ACTION: Users interested in a wide-field NIR camera at APO should
send a paragraph or two to their Users Committee representative with
copies to Suzanne and Bruce with an expression of interest and/or
questions regarding this proposed instrument.
STATUS: Closed, see above discussion.
ACTION: We need each institution's Users Committee member to poll
their users and send science highlights (since Oct. '06) to Suzanne
and Bruce, not later than 15 October, for inclusion in the annual
report to the ARC Board.
STATUS: Closed, completed.
ACTION: The decision on whether or not to proceed with moving forward
on the wide-field NIR camera proposal will be deferred to the November
Users Committee meeting. In the meantime, committee members should
make one more pass at soliciting expressions of interest from their
user communities. Suzanne and Bruce will apprise the instrument PI on
where we are.
STATUS: Closed, see above discussion.
ACTION: Committee members to look at the new user webpages, and
report back to Jon Holtzman and Gabrelle Saurage suggestions for
content and organization.
STATUS: Open, see above discussion
[new actions]:
ACTION: Jon Bally will write a section with figures on Fowler
sampling for inclusion in the NIC-FPS user documentation.
**********************************
Next meeting:
The next Users Committee phonecon will be on Monday, 10 December, at
8:30 AM Pacific Time. The agenda and other materials will be sent to
the committee members during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1073 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO